# Bastani, Bastani, and McLaughlin

### v2

@@ 2.2 Approach
To characterize $\Pi_{\Lambda}$ we need to solve this convex problem for all values of $\lambda\in[0,\tau_{max}]$. The first order condition of this optimization problem does not contain $\lambda$, instead giving a form for $\pi_{n,t}$ in terms of dual variables whose optimal values are induced by $\lambda$. The Lagrangian associated with all problems forming $\Pi_{\Lambda}$ is,
$\mathcal{L}(\pi,\xi,\beta,\kappa)=s^{2}(\pi)+\xi\left(\lambda-\tau(\pi)\right)+\sum_{n=1}^{N}\beta_{n}\left(1-\sum_{t=0}^{K}\pi_{n,t}\right)-\sum_{n=1}^{N}\sum_{t=0}^{K}\kappa_{n,t}\pi_{n,t},$
@@
Thus, the first order condition with respect to $\pi_{n,t}^{*}$ for an optimal solution $(\pi^{*},\xi^{*},\beta^{*},\kappa^{*})$ is
$\frac{2}{N^{2}}\left[\frac{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}{\pi^{o}_{n,t}}\left(\pi_{n,t}^{*}-\pi^{o}_{n,t}\right)-\mu_{n,t}\sum_{v=0}^{K}\mu_{n,v}\left(\pi_{n,v}^{*}-\pi^{o}_{n,v}\right)\right]+\frac{\xi^{*}}{N}\mu_{n,t}-\beta_{n}^{*}-\kappa_{n,t}^{*}=0.$
We can eliminate $\beta^{*}_{n}$ using the primal constraint $\sum_{t=0}^{K}\pi_{n,t}^{*}=1$ and then reorganize this system of equations to solve for $\pi^{*}$ in terms of $\xi^{*}$ and $\kappa^{*}$, {repeatproposition} We have
$\pi^{*}_{n,t}$ $=\pi^{o}_{n,t}\left(1+\frac{Q_{n}^{-1}\left(\xi^{*}+N\sum_{v=0}^{K}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}\kappa^{*}_{n,v}\left(\mu_{n,v}-\tilde{\mu}_{n}\right)\right)\left(\mu_{n,t}-\tilde{\mu}_{n}\right)+N\left(\kappa^{*}_{n,t}-\tilde{\kappa}^{*}_{n}\right)}{2\left(\mu_{n,t}^{2}+\sigma_{n,t}^{2}\right)}N\right)$
@@
To convert Proposition 2.2 into a characterization of $\Pi_{\Lambda}$, we solve for the values of all dual variables except $\xi^{*}$, leaving a one-dimensional parameterization of $\Pi_{\Lambda}$ in terms of $\xi^{*}$.
2.3 Binary Characterization
@@
For intuition, we first characterize $\Pi_{\Lambda}$ for the case of a binary treatment $(K=1)$ before describing the general solution. The first order condition given in Proposition 2.2 is independent across units, thus, for any $\xi^{*}$ we can calculate $\pi_{n}^{*}(\xi^{*})$ in terms of $\pi_{n}^{o}$, $\mu_{n}$, and $\sigma^{2}_{n}$. This provides insight into how the dual variable determines optimal solution propensities in terms of the relative quality of the treatment for each unit.
In case of binary treatments ($K=1$), the expectation maximizing treatment for each unit is given by the sign of its conditional average treatment effect (CATE). If the variance of the conditional treatment effect for unit $n$ is large relative to the CATE, there is a reasonable chance IPW will view the expectation maximizing treatment as a loss, introducing a large amount of variance to its evaluation. To counteract this, solutions along $\Pi_{\Lambda}$ move from the observational policy (where there is no evaluation variance) towards the expectation maximizing policy, with units whose CATE is large relative to the variance of their conditional treatment effect moving the most.
We characterize the policies $\pi\in\Pi_{\Lambda}$ in terms of $\pi_{n,1}$, the propensity to give the treatment to unit $n$, noting that $\pi_{n,0}=1-\pi_{n,1}$. {repeattheorem} For $K=1$ (binary treatment), we have
$\Pi_{\Lambda}$ $=\bigcup_{\xi\geq 0}\left\{\pi^{*}(\xi)\right\}$ where $\pi^{*}_{n,1}(\xi)$ $=\min\left(\left[\pi^{o}_{n,1}+\frac{N\xi\left(\mu_{n,1}-\mu_{n,0}\right)}{2\left(\frac{\mu^{2}_{n,0}+\sigma^{2}_{n,0}}{1-\pi^{o}_{n,1}}+\frac{\mu^{2}_{n,1}+\sigma^{2}_{n,1}}{\pi_{n,1}^{o}}-\left(\mu_{n,1}-\mu_{n,0}\right)^{2}\right)}\right]^{+},1\right),$
and where $\cdot]^{+}=\max(0,\cdot)$. In other words, the policies in $\Pi_{\Lambda}$ are parametrized by the optimal value of the dual variable $\xi$. As $\xi$ increases $\pi^{*}(\xi)$ moves linearly towards the expectation maximizing policy at a rate that increases with the size of the observational dataset. Each unit’s policy is determined independent of one another, thus when one unit’s treatment matches the expectation maximizing treatment, it remains there while other units’ treatments continue to move with $\xi$. Above some $\xi_{\max}$, $\pi^{*}(\xi)$ simply returns to the expectation maximizing policy.
Each unit’s treatment propensity changes in proportion to the IPW estimator’s evaluation of the conditional average treatment effect over the variance of the conditional treatment effect. The IPW estimator upweights all outcomes by their inverse propensity $Y^{o}_{n}/\pi^{o}_{n,T_{n}^{o}}$, and then uses each weighted outcome according to $\pi_{n,T_{n}^{o}}$ to estimate the impact of $\pi$. Thus, according to IPW, the mean and variance of the conditional treatment effect are
$\mathbb{E}\left[\left(\mathbf{1}[T_{n}^{o}=1]-\mathbf{1}[T_{n}^{o}=0]\right)\frac{Y_{n}^{o}}{\pi^{o}_{n,T_{n}^{o}}}\right]$ $=\mu_{n,1}-\mu_{n,0},\hskip 20.0pt\text{and}$
$\mathrm{Var}\left[\left(\mathbf{1}[T_{n}^{o}=1]-\mathbf{1}[T_{n}^{o}=0]\right)\frac{Y_{n}^{o}}{\pi^{o}_{n,T_{n}^{o}}}\right]$ $=\frac{\mu_{n,0}^{2}+\sigma_{n,0}^{2}}{1-\pi^{o}_{n,1}}+\frac{\mu_{n,1}^{2}+\sigma_{n,1}^{2}}{\pi_{n,1}^{o}}-\left(\mu_{n,1}-\mu_{n,0}\right)^{2},$
respectively. As $\xi$ increases, each unit’s treatment propensity $\pi^{*}(\xi)$ changes (if possible) in terms of the ratio of the CATE to the variance of the conditional treatment effect as estimated by IPW.
2.4 Multiarm Characterization

### v3

@@ 2.2 Convex Relaxation Framework
To characterize $\Pi_{\Lambda}$ we need to solve this convex problem for all values of $\lambda\in[0,\tau_{\text{max}}]$. The first order condition of this optimization problem does not contain $\lambda$, instead giving a form for $\pi_{n,t}$ in terms of dual variables whose optimal values are induced by $\lambda$. The Lagrangian associated with all problems forming $\Pi_{\Lambda}$ is,
$\mathcal{L}(\pi,\zeta,\beta,\kappa)=s^{2}(\pi)+\zeta\left(\lambda-\tau(\pi)\right)+\sum_{n=1}^{N}\beta_{n}\left(1-\sum_{t=0}^{K}\pi_{n,t}\right)-\sum_{n=1}^{N}\sum_{t=0}^{K}\kappa_{n,t}\pi_{n,t},$
@@
Thus, the first order condition with respect to $\pi_{n,t}^{*}$ for an optimal solution $(\pi^{*},\zeta^{*},\beta^{*},\kappa^{*})$ is
$\frac{2}{N^{2}}\left[\frac{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}{\pi^{o}_{n,t}}\left(\pi_{n,t}^{*}-\pi^{o}_{n,t}\right)-\mu_{n,t}\sum_{v=0}^{K}\mu_{n,v}\left(\pi_{n,v}^{*}-\pi^{o}_{n,v}\right)\right]+\frac{\zeta^{*}}{N}\mu_{n,t}-\beta_{n}^{*}-\kappa_{n,t}^{*}=0.$
We can eliminate $\beta^{*}_{n}$ using the primal constraint $\sum_{t=0}^{K}\pi_{n,t}^{*}=1$ and then reorganize this system of equations to solve for $\pi^{*}$ in terms of $\zeta^{*}$ and $\kappa^{*}$, {repeatproposition} We have
$\pi^{*}_{n,t}$ $=\pi^{o}_{n,t}\left(1+\frac{Q_{n}^{-1}\left(\zeta^{*}+N\sum_{v=0}^{K}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}\kappa^{*}_{n,v}\left(\mu_{n,v}-\tilde{\mu}_{n}\right)\right)\left(\mu_{n,t}-\tilde{\mu}_{n}\right)+N\left(\kappa^{*}_{n,t}-\tilde{\kappa}^{*}_{n}\right)}{2\left(\mu_{n,t}^{2}+\sigma_{n,t}^{2}\right)}N\right)$
@@
To convert Proposition 2.2 into a characterization of $\Pi_{\Lambda}$, we solve for the values of all dual variables except $\zeta^{*}$, leaving a one-dimensional parameterization of $\Pi_{\Lambda}$ in terms of $\zeta^{*}$.
3 Single Treatment Setting
For intuition, we first consider for the case of comparing a single treatment to a control (i.e., $K=1$). We begin by solving our convex relaxation to obtain a superset of the Pareto frontier (Section 3.1). Then, we characterize the structure of the Pareto frontier based on this result (Section 3.2). Finally, we discuss implications of our characterization for policy learning methodologies (Section 3.3). Section 4 generalizes our characterization results to the multi-treatment setting.
3.1 Superset of the Pareto Frontier
@@
We begin by characterizing the Pareto frontier $\Pi_{\Lambda}$. The first order condition given in Proposition 2.2 is independent across units, thus, for any $\zeta^{*}$ we can calculate $\pi_{n}^{*}(\zeta^{*})$ in terms of $\pi_{n}^{o}$, $\mu_{n}$, and $\sigma^{2}_{n}$. This provides insight into how the dual variable determines optimal solution propensities in terms of the relative quality of the treatment for each unit. We characterize the policies $\pi\in\Pi_{\Lambda}$ in terms of $\pi_{n,1}$, the propensity to give the treatment to unit $n$, noting that $\pi_{n,0}=1-\pi_{n,1}$. For convenience, define $\tau(\zeta)=\tau(\pi^{*}(\zeta))$, $s^{2}(\zeta)=s^{2}(\pi^{*}(\zeta))$, and $Z(\zeta)=\mathbb{E}[Z(\pi^{*}(\zeta))]=\tau(\zeta)/s(\zeta)$. A proof for Theorem 3.1 is given in Appendix 8. {repeattheorem} For the single-treatment setting (i.e., $K=1$), we have $\Pi_{\Lambda}=\bigcup_{\zeta\geq 0}\{\pi^{*}(\zeta)\}$. Also,
$\pi^{*}_{n,t}(\zeta)=\pi_{n,t}^{o}+\psi_{t}\alpha_{n}\qquad\qquad\tau(\zeta)=\frac{1}{N}\sum_{n=1}^{N}\tau_{n}\alpha_{n}\qquad\qquad s^{2}(\zeta)=\frac{1}{N^{2}}\sum_{n=1}^{N}\eta_{n}\alpha_{n}^{2},$ (1)
where
$\psi_{t}=\begin{cases}1&\text{if }t=1\\ -1&\text{if }t=0\end{cases}\qquad\qquad\alpha_{n}=\begin{cases}\min\left\{\dfrac{N\zeta\tau_{n}}{2\eta_{n}},1-\pi_{n,1}^{o}\right\}&\text{if }\tau_{n}>0\\ 0&\text{if }\tau_{n}=0\\ \max\left\{\dfrac{N\zeta\tau_{n}}{2\eta_{n}},-\pi_{n,1}^{o}\right\}&\text{if }\tau_{n}<0,\end{cases}$
and where
$\tau_{n}=\mu_{n,1}-\mu_{n,0}\qquad\qquad\qquad\eta_{n}=\sum_{t=0}^{1}\frac{\sigma^{2}_{n,t}}{\pi^{o}_{n,t}}+\left(\sum_{t=0}^{1}\sqrt{\frac{1-\pi^{o}_{n,t}}{\pi^{o}_{n,t}}}\mu_{n,t}\right)^{2}.$
Here, $\tau_{n}$ is the treatment effect for unit $n$, and $\eta_{n}$ is the variance of its IPW estimate: [^3] The second term of $\eta_{n,t}$ involving $\mu_{n,t}$ may seem strange; it arises because the IPW estimator is not invariant to additive constants to the outcome variable. Indeed, normalizing outcomes can reduce variance of IPW, resulting in the augmented IPW (AIPW) estimator (Robins et al. 1994).
$\tau_{n}$ $=\mathbb{E}\left[\left(\mathbbm{1}(T_{n}^{o}=1)-\mathbbm{1}(T_{n}^{o}=0)\right)\frac{Y_{n}^{o}}{\pi^{o}_{n,T_{n}^{o}}}\right]$
$\eta_{n}$ $=\text{Var}\left[\left(\mathbbm{1}(T_{n}^{o}=1)-\mathbbm{1}(T_{n}^{o}=0)\right)\frac{Y_{n}^{o}}{\pi^{o}_{n,T_{n}^{o}}}\right]$
Furthermore, $\Pi_{\Lambda}$ is parameterized by the optimal value of the dual variable $\zeta$. It starts from the observational policy $\pi^{*}(0)=\pi^{o}$. As $\zeta$ increases, the treatment probability increases or decreases (depending on the sign of $\tau_{n}$) continuously at a rate of $N\tau_{n}/(2\eta_{n})$ until it binds to either zero or one; the binding order is determined by the magnitude of this rate. This rate prioritizes units with higher expected treatment effect and lower estimation variance; notably, it is independent of the other units except through the total number of units $N$. The binding value optimizes the expected treatment effect; also, letting $\zeta_{\text{max}}$ denote the smallest $\zeta$ for which which the treatment probabilities are bound for all units, then $\pi^{*}(\zeta)$ is the expectation maximizing policy for all $\zeta\geq\zeta_{\text{max}}$.
3.2 Structure of the Pareto Frontier
Figure 1: Structure of the Pareto frontier.
Theorem 3.1 provides detailed insights into the structure of the Pareto frontier. First, we define
$\mathcal{N}_{u}(\zeta)=\left\{n\in[N]\biggm|\alpha_{n}=\frac{N\zeta\tau_{n}}{\eta_{n}}\right\},$
i.e., $\mathcal{N}_{u}(\zeta)$ is the set of units for which $\pi_{n,t}^{*}(\zeta)$ is not bound to zero or one (but including the boundary case); we denote its complement by $\mathcal{N}_{b}(\zeta)=[N]\setminus\mathcal{N}_{u}(\zeta)$.
{repeatproposition}
The treatment effect $\tau(\zeta)$ is monotonically increasing in $\zeta$, and this increase is strict for all $\zeta>0$ such that $|\mathcal{N}_{u}(\zeta)|<N$.
Proof.
Proof. This result follows immediately from Eq. (1). ∎∎
The behavior of the $z$-score is only slightly more complicated (though its proof is less trivial): {repeatproposition} The $z$-score $Z(\zeta)$ is monotonically decreasing in $\zeta$, and this decrease is strict for all $\zeta>0$ such that $1\leq|\mathcal{N}_{u}(\zeta)|<N$; furthermore, it is constant for $\zeta>0$ such that $|\mathcal{N}_{u}(\zeta)|=0$. Before we provide a proof, we first describe the implications of Proposition 3.21 for the structure of the Pareto frontier, which is illustrated in Figure 1. First, note that the observational policy is $\pi^{o}=\pi^{*}(0)$; thus, we have $\tau(0)=0$ (i.e., observational policy has zero improvement over itself) and $Z(0)$ is undefined since $s^{2}(0)=0$ (intuitively, we can never “prove” that the observational policy is better than itself). Next, define
$\zeta_{\text{min}}=\inf\{\zeta\geq 0\mid\mathcal{N}_{b}(\zeta)\neq\varnothing\}\qquad\qquad\qquad\zeta_{\text{max}}$ $=\inf\{\zeta\geq 0\mid\mathcal{N}_{u}(\zeta)=\varnothing\}.$
Intuitively, $\zeta_{\text{min}}$ is the smallest value of $\zeta$ for which some unit has treatment probability bound to zero or one; analogously, $\zeta_{\text{max}}$ is the smallest value at which all units become bound to zero or one (so $\pi^{*}(\zeta_{\text{max}})\in\operatorname*{\arg\max}_{\pi}\tau(\pi)$ is the maximizer of the expected treatment effect).
By Proposition 1, the $z$-score is constant for $\zeta\in(0,\zeta_{\text{min}}]$, and by Proposition 3.2, the expected treatment effect is strictly increasing along this interval. Thus, $\zeta_{\text{min}}$ is a strictly better choice than the remainder of the interval, so we prune $(0,\zeta_{\text{min}})$ from the frontier. Now, Propositions 3.21 say that on $\zeta_{\text{min}},\zeta_{\text{max}}]$, the $z$-score is strictly decreasing and the expected treatment effect is strictly increasing; thus, all policies $\pi^{*}(\zeta)$ for $\zeta_{\text{min}}\leq\zeta\leq\zeta_{\text{max}}$ are part of the frontier. Finally, the policy becomes constant for $\zeta\geq\zeta_{\text{max}}$, so we also prune $(\zeta_{\text{max}},\infty)$ from the frontier. Thus: {repeatcorollary} The Pareto frontier is $\Pi_{\text{opt}}=\{\pi^{*}(\zeta)\mid\zeta\in[\zeta_{\text{min}},\zeta_{\text{max}}]\}$.
Proof.
Proof of Proposition 1. Note that
$Z^{\prime}(\zeta)=\frac{\tau^{\prime}(\zeta)s(\zeta)-s^{\prime}(\zeta)\tau(\zeta)}{s^{2}(\zeta)}.$
Thus, it suffices to show that $\frac{s^{\prime}(\zeta)}{s(\zeta)}-\frac{\tau^{\prime}(\zeta)}{\tau(\zeta)}\geq 0$, and that this inequality is strict when $|\mathcal{N}_{b}(\zeta)|>0$. Note that $(s^{2})^{\prime}(\zeta)=2s(\zeta)s^{\prime}(\zeta)$, so in fact it suffices to show that
$\Delta(\zeta)\coloneqq\frac{(s^{2})^{\prime}(\zeta)}{2s^{2}(\zeta)}-\frac{\tau^{\prime}(\zeta)}{\tau(\zeta)}\geq 0.$
By our definition of $\mathcal{N}_{u}(\zeta)$ and $\mathcal{N}_{b}(\zeta)$, we have
$\tau(\zeta)$ $=\frac{1}{N}\sum_{n\in\mathcal{N}_{b}(\zeta)}\tau_{n}\alpha_{n}+\frac{\zeta}{2}\sum_{n\in\mathcal{N}_{u}(\zeta)}\frac{\tau_{n}^{2}}{\eta_{n}}$
$s^{2}(\zeta)$ $=\frac{1}{N^{2}}\sum_{n\in\mathcal{N}_{b}(\zeta)}\eta_{n}\alpha_{n}^{2}+\frac{\zeta^{2}}{4}\sum_{n\in\mathcal{N}_{u}(\zeta)}\frac{\tau_{n}^{2}}{\eta_{n}}.$
Defining $A(\zeta)=N^{-1}\sum_{n\in\mathcal{N}_{b}(\zeta)}\tau_{n}\alpha_{n}$, $B(\zeta)=N^{-2}\sum_{n\in\mathcal{N}_{b}(\zeta)}\eta_{n}\alpha_{n}^{2}$, and $D(\zeta)=\sum_{n\in\mathcal{N}_{u}(\zeta)}(\tau_{n}^{2}/\eta_{n})$, then
$\tau(\zeta)=A(\zeta)+\frac{\zeta}{2}D(\zeta),\qquad s^{2}(\zeta)=B(\zeta)+\frac{\zeta^{2}}{4}D(\zeta).$
Since $A(\zeta)$, $B(\zeta)$, and $D(\zeta)$ are piecewise constant, we have
$\Delta(\zeta)$ $=\frac{D(\zeta)}{(4B(\zeta)/\zeta)+\zeta D(\zeta)}-\frac{D(\zeta)}{2A(\zeta)+\zeta D(\zeta)}.$
Thus, it suffices to show that $2B(\zeta)/\zeta\geq A(\zeta)$. Note that by definition, for any $n\in\mathcal{N}_{b}$, we have
$|\alpha_{n}|>\frac{N\zeta|\tau_{n}|}{2\eta_{n}}$
As a consequence, we have
$\frac{2B(\zeta)}{\zeta}=\frac{1}{N}\sum_{n\in\mathcal{N}_{b}(\zeta)}\frac{2\eta_{n}\alpha_{n}^{2}}{N\zeta}\geq\frac{1}{N}\sum_{n\in\mathcal{N}_{b}(\zeta)}|\tau_{n}|\cdot|\alpha_{n}|=A(\zeta),$
where the inequality is strict if $\mathcal{N}_{b}(\zeta)\neq\varnothing$; here, we have used the fact that $|\tau_{n}|\cdot|\alpha_{n}|=\tau_{n}\alpha_{n}$
So far, we have ignored the fact that $Z(\zeta)$ is not everywhere differentiable. However, it is everywhere continuous and differentiable except potentially on a finite set of points. As a consequence, if $Z^{\prime}(\zeta)<0$ on an open interval $(\zeta_{\text{min}},\zeta_{\text{max}})$ except at points where $Z^{\prime}$ is not defined, then it follows that $Z$ is strictly monotone everywhere on that interval.
It remains to show that $Z(\zeta)$ is constant when $\mathcal{N}_{b}(\zeta)=\varnothing$. In this case, we have
$Z(\zeta)=\frac{\tau(\zeta)}{s(\zeta)}=\frac{\frac{\zeta}{2}\sum_{n=1}^{N}\frac{\tau_{n}^{2}}{\eta_{n}}}{\sqrt{\frac{\zeta^{2}}{4}\sum_{n=1}^{N}\frac{\tau_{n}^{2}}{\eta_{n}}}}=\sqrt{\sum_{n=1}^{N}\frac{\tau_{n}^{2}}{\eta_{n}}},$
which does not depend on $\zeta$. The claim follows. ∎∎
3.3 Discussion
Here, we discuss several implications of our results for policy learning algorithms targeting statistically significant performance improvements. While our discussion is based on the characterization in a specialized setting, we show that our key insights extend to the general setting in Section 4; thus, our discussion considers the general case.
Comparing policy learning methodologies.
Prior work has proposed ad hoc techniques for policy learning targeting statistical significance. They can be grouped into three strategies:
  • Variance regularization: These approaches select policies that minimize some combination of the mean and variance of the treatment effect. In the idealized setting where all $\mu_{n,t}$ and $\sigma_{n,t}$ are known, variance regularization computes a policy along our Pareto frontier.
  • Treatment aggregation: These approaches merges similar treatments, which mitigates the winner’s curse since they select among fewer candidate policies. Note that this strategy can only be applied when multiple treatments are available.
  • Unit aggregation: These approaches aggregate similar units; with the same goal of reducing the number of candidate policies as treatment aggregation.
To compare these approaches, we introduce the following standard definitions.
Definition 3.1
$\pi$ is deterministic if $\forall n\in[N],t\in\mathcal{T}$, $\pi_{n,t}\in\{0,1\}$; otherwise, it is stochastic.
Definition 3.2
$\pi$ is untargeted if $\forall t\in\mathcal{T}$, $\pi_{n,t}=\pi_{t}$ is independent of $n$; otherwise, it is targeted.
One of our key insights is that policies along the Pareto frontier are largely stochastic. {repeatcorollary} All policies $\pi\in\Pi_{\text{opt}}$ are stochastic except $\pi^{*}=\pi^{*}(\zeta_{\text{max}})$. Now, we consider the effect of treatment aggregation on the policy space. Treatment aggregation typically considers a large number of treatment options, but does not consider targeting—i.e., we have a finite policy class $\Pi=\{\pi^{t}\}_{t\in\mathcal{T}}$, where $\pi^{t}_{n,t^{\prime}}=\mathbbm{1}(t^{\prime}=t)$ [^4] Targeting can be in principle encoded by having the “treatment” vary depending on the unit, but this is not typical.. In other words, policy $\pi^{t}$ deterministically assigns treatment $t$ to all units. Now, merging treatments $t_{1},...,t_{H}\in\mathcal{T}$ is mathematically achieved by creating a new policy $\tilde{\pi}$ that uniformly randomly selects among these treatments—i.e.,
$\tilde{\pi}_{n,t}=\frac{1}{H}\sum_{h=1}^{H}\pi_{n,t}^{t_{h}}.$
Then, $\pi^{t_{1}},...,\pi^{t_{H}}$ are removed from $\Pi$ and $\tilde{\pi}$ is added to $\Pi$. Note that $\tilde{\pi}$ is stochastic, which is consistent with our insight that Pareto optimal policies are stochastic. However, in our analysis in Section 4, we show that the policy $\pi^{*}(\zeta_{\text{min}})$ is achieved by increasing $\zeta$ until $\pi_{n,t}^{*}(\zeta)$ binds to zero for any unit $n\in[N]$ and treatment $t\in\mathcal{T}$; thus, $\pi^{*}(\zeta_{\text{min}})$ typically randomizes over all treatments. It may be possible to improve treatment aggregation by randomly selecting among all treatments in a non-uniform way rather than merging a few high-value treatments.
Unit aggregation approaches also assume a finite policy class $\Pi$, where in this case each policy $\pi\in\Pi$ is a targeted policy; then, it identifies similar units $n_{1},...,n_{H}\in[N]$ and imposes the constraint that $\pi_{n_{h}}=\pi_{n_{h^{\prime}}}$ for all $h,h^{\prime}\in[H]$. A typical strategy is to simply remove policies that do not satisfy this constraint from $\Pi$; thus, if $\Pi$ initially only contains deterministic policies, then unit aggregation never introduces any stochasticity into $\Pi$. As a consequence, we might expect these approaches to be more effective if $\Pi$ initially contains stochastic policies.
Statistical power.
So far, we have been assuming the observational dataset is fixed; however, a common question in practice is how to collect the sample $\mathcal{O}$ to achieve a target level of significance. To address this question, we need to understand $Z(\zeta_{\text{min}})$, which is the best possible $z$-score that can be achieved along the Pareto frontier. In fact, we have an explicit formula for $Z(\zeta_{\text{min}})$: {repeatcorollary} The best possible $z$-score is achieved by $\pi^{*}(\zeta_{\text{min}})$, and equals
$Z(\zeta_{\text{min}})=\sqrt{\sum_{n=1}^{N}Z_{n}},\qquad\text{where}\qquad Z_{n}=\frac{\tau_{n}^{2}}{\eta_{n}}.$
This result follows from the proof of Proposition 1. Given assumptions about $\mu_{n,0}$, $\mu_{n,1}$, and $\sigma_{n}$, it can be used to calculate the number of samples needed for a target level of statistical significance. It also quantifies the value of each unit toward statistical significance, which might help identify promising test subjects when recruitment is costly (Kouvelis et al. 2017, Tian et al. 2023). For instance, suppose we further assume the following: (1) $\pi^{o}$ is an RCT policy (i.e., $\pi^{o}_{n,t}=1/2$ for all $n\in[N]$ and $t\in\mathcal{T}$), (2) the outcome variances are independent of treatment (i.e., $\sigma_{n,t}=\sigma_{n}$ for some $\sigma_{n}$), and (3) outcomes are centered (i.e., $\mu_{n,1}=-\mu_{n,0}=\mu_{n}$ for some $\mu_{n}$; for instance, the augmented IPW estimator aims to center outcomes). Then, we have
$Z_{n}=\frac{\mu_{n}^{2}}{\sigma_{n}^{2}},$
which is exactly the signal-to-noise ratio (SNR) for unit $n$. For instance, note that
$\frac{\partial}{\partial Z_{n}}Z(\zeta_{\text{min}})=\frac{1}{2Z(\zeta_{\text{min}})}.$
From this result, we observe that the marginal value of increasing the SNR $Z_{n}$ of unit $n$ only depends on the current $z$-score $Z(\zeta_{\text{min}})$. Intuitively, the marginal value decreases proportional to $Z(\zeta_{\text{min}})$. Thus, it is always better to allocate resources to recruiting units with higher SNR, but the overall efficacy of increasing SNR has diminishing returns.
Finally, we can compare $\pi^{*}(\zeta_{\text{min}})$ to $\pi^{*}(\zeta_{\text{max}})$, which is given by
$\pi^{*}_{n,t}(\zeta_{\text{max}})$ $=\begin{cases}1&\text{if }\tau_{n}>0\\ \pi^{o}_{n,t}&\text{if }\tau_{n}=0\\ 0&\text{if }\tau_{n}<0.\end{cases}$
While $Z(\zeta_{\text{max}})$ can be complicated in general, it simplifies when $\pi^{o}$ is uniform. {repeatcorollary} Suppose that $\pi_{n,t}^{o}=1/2$ for all $n\in[N]$ and $t\in\mathcal{T}$. Then, we have
$Z(\zeta_{\text{max}})=\frac{\sum_{n=1}^{N}|\tau_{n}|}{\sqrt{\sum_{n=1}^{n}\eta_{n}}}.$
This result follows from the proof of Proposition 1. Under conditions (1), (2), and (3) above,
$Z(\zeta_{\text{max}})=\frac{\sum_{n=1}^{N}|\mu_{n}|}{\sqrt{\sum_{n=1}^{N}\sigma_{n}^{2}}}.$
In particular, in this case, it is no longer true that $Z(\zeta_{\text{max}})$ can be expressed as in terms of the SNRs of individual units. In general, $Z(\zeta_{\text{min}})\geq Z(\zeta_{\text{max}})$; this inequality follows from Proposition 1, but it also follows directly from Sedrakyan’s inequality (Sedrakyan and Sedrakyan 2018). It is tight when $\mu_{n}=\mu$ and $\sigma_{n}=\sigma$ are constant; indeed, in this case, it is easy to see that $\pi^{*}(\zeta_{\text{min}})=\pi^{*}(\zeta_{\text{max}})$. To understand when there is a gap, consider the following:
  • Arbitrarily large gap: Consider two units with $\mu_{1}=\mu_{2}=1$, $\sigma_{1}=1$, and $\sigma_{2}=\sigma$; then
    $Z(\zeta_{\text{min}})=\sqrt{1+\frac{1}{\sigma^{2}}}\to\infty\qquad\qquad Z(\zeta_{\text{max}})=\frac{2}{\sqrt{1+\sigma^{2}}}\to 2$
    as $\sigma\to 0$. Intuitively, $\pi^{*}(\zeta_{\text{min}})$ “drops” the first unit from the analysis (i.e., taking $\pi_{1,t}^{*}(\zeta_{\text{min}})\to\pi_{1,t}^{o}$ while $\pi_{2,t}^{*}(\zeta_{\text{min}})\to\mathbbm{1}(t=1)$), so as $\sigma\to 0$, it achieves complete confidence in policy improvement. In contrast, $\pi^{*}(\zeta_{\text{max}})$ always treats both units (i.e., taking $\pi_{n,t}^{*}(\zeta_{\text{max}})=\mathbbm{1}(t=1)$; thus, it always retains the variance from treating the first unit, bounding its confidence.
  • Asymptotic gap: There is a gap between the two even as $N\to\infty$. We can easily show a constant multiplicative gap; consider two unit types, where there are $N$ units of the first type (with $\mu_{n}=1$ and $\sigma_{n}=1$) and $N$ of the second type (with $\mu_{n}=1$ and $\sigma_{n}=\sigma$, for $\sigma\ll 1$). Then, we have
    $Z(\zeta_{\text{min}})=\sqrt{N}\cdot\sqrt{1+\frac{1}{\sigma^{2}}}\geq\sqrt{N}\cdot\frac{1}{\sigma}\qquad\qquad Z(\zeta_{\text{max}})=\sqrt{N}\cdot\frac{2}{\sqrt{1+\sigma^{2}}}\leq\sqrt{N}\cdot 2$
    We can also show a non-constant gap, though it requires the variance of each unit to grow with $n$; specifically, letting $\mu_{n}=1$ for all $n$ and $\sigma_{n}=\sqrt{n}$, then we have
    $Z(\zeta_{\text{min}})=\sqrt{\sum_{n=1}^{N}\frac{1}{n}}\geq\sqrt{\log N}\qquad\qquad Z(\zeta_{\text{max}})=\frac{N}{\sqrt{\sum_{n=1}^{N}n}}=\frac{\sqrt{2}N}{\sqrt{N(N+1)}}\leq\sqrt{2}.$
    Intuitively, this scenario involves adding higher and higher variance units to the sample, which gradually improves $Z(\zeta_{\text{min}})$ but not $Z(\zeta_{\text{max}})$.
  • Non-monotonicity of $Z(\zeta_{\text{max}})$: Consider two unit types, where there are $N$ units of the first type (with $\mu_{n}=\mu\ll 1$ and $\sigma_{n}=1$) and $M\gg 1$ of the second (with $\mu_{n}=1$ and $\sigma_{n}=1$). Then
    $Z(\zeta_{\text{min}})=\sqrt{N\mu+M}\qquad\qquad Z(\zeta_{\text{max}})=\frac{N\mu+M}{\sqrt{N+M}}.$
    While $Z(\zeta_{\text{min}})$ is monotonically increasing in $N$ (in general, it is monotonically increasing as units are added), this is not true for $Z(\zeta_{\text{max}})$ for sufficiently small $\mu$ and large $M$. Intuitively, $\pi^{*}(\zeta_{\text{max}})$ initially performs well due to the $M$ “good” units, so adding a small number of “bad” units actually reduces its $z$-score by adding variance. Eventually, adding sufficiently many bad units overcomes this variance to improve the $z$-score. If we choose $\sigma_{n}$ to increase with $n$, $Z(\zeta_{\text{max}})$ can be made to monotonically decrease in $n$; specifically, letting $\mu_{n}=1$ and $\sigma_{n}=n$ for all $n$, then we have
    $Z(\zeta_{\text{min}})=\sqrt{\sum_{n=1}^{N}\frac{1}{n^{2}}}\qquad\qquad Z(\zeta_{\text{max}})=\frac{N}{\sqrt{\sum_{n=1}^{N}n^{2}}}=\frac{N}{\sqrt{N(N+1)(2N+1)/6}}.$
    As $N\to\infty$, then $Z(\zeta_{\text{max}})\to 0$ and $Z(\zeta_{\text{min}})\to\sqrt{\pi^{2}/6}$ (here, $\pi$ refers to the number, not the policy). That is, if we add units with successively higher variance by a sufficient increment, this variance actually decreases the $z$-score of $\pi^{*}(\zeta_{\text{max}})$.
4 Multiple Treatment Setting
Next, we consider the general case of multiple treatments. Mirroring our analysis in the single-treatment analysis, we begin by computing the solutions to our optimization problem, which form a superset of the Pareto frontier (Section 4.1). Then, we establish the structure of the Pareto frontier, which is essentially identical to the structure in Figure 1 for the single-treatment setting (Section 4.2). Finally, we provide results on the structure of policies along the Pareto frontier (Section 4.3) and the best possible statistical power (Section 4.4). Proofs for all claims are given in Appendix 8.
4.1 Superset of the Pareto Frontier